### v1

@@ 4.1 Test-Time Distribution Sizes
EasySpec 1.61($\uparrow$1.57x) 1.57 3.24(3.64x) 0.87 1.48($\uparrow$1.52x) 1.54 3.02(3.97x) 0.91
@@
4.2 Main Results
4.2.1 Effectiveness
To illustrate the effectiveness of EasySpec, we conducted experiments on Llama-3-70B-Instruct and Qwen2-72B-Instruct across datasets, with temperatures=0 and 0.8. The results are shown in Table 2. We measure the running time per 100 tokens of the drafting (d) and verification (v) stages and the total process (total). $\alpha$ represents the acceptance rates.
The results show that EasySpec consistently accelerates the overall execution of speculative decoding across all datasets. As the verification stage (the large models) can be accelerated by TP, the drafting stage becomes the bottleneck, accounting for nearly 2/3 of the total running time. EasySpec accelerates the drafting stage by up to 1.62x, sometimes surpassing the acceleration rates of TP for the large models (around 1.5x), meaning that the optimized drafting-stage multi-GPU utilization is as much efficient as TP can achieve on the verification stage. Meanwhile, the drop of speculation accuracy is no more than 7%, suggesting that fuzzy speculation is sufficiently accurate in most cases.
We also evaluated EasySpec on task-specific models: Qwen2-Math-72B-Instruct and Qwen2.5-Coder-32B-Instruct. Table 3 shows the results of acceleration on the specific-task datasets MATH and HumanEval. On these datasets, the original speculation results of the draft models is already highly accurate (with the acceptance rates greater than 97%), indicating that the speculation is inherently simple. As a result, fuzzy speculation incurs virtually no drop in accuracies, achieving an ideal drafting-stage acceleration with little-to-no cost.
Table 3: Running time and acceptance rates of task-specific models on datasets of their corresponding tasks.
Q2.5-Coder-32B(7B)-Instruct
Datasets Methods d v total $\alpha$
Human Eval vanilla - - 7.13 -
+tree 2.15 1.23 3.38 0.97
EasySpec 1.43 1.23 2.66 0.97
-0.72 +0.00 -0.72 -0.00
Q2-Math-72B(7B)-Instruct
MATH vanilla - - 11.55 -
+tree 2.05 1.51 3.55 0.99
EasySpec 1.40 1.52 2.91 0.98
-0.65 +0.01 -0.64 -0.01
4.2.2 Generalization and Stability
An intermediate question is that whether we can use an even smaller model to reduce drafting latency. Firstly, such a model often does not exist within the same series, while training a smaller draft model is non-trivial and the trained model often lacks the ability of generalization across base models and datasets Fu et al. . Secondly, an overly smaller model tends to exhibit unstable speculation performance, while models in the same series inherently have an alignment in prediction behaviors due to shared tokenizer, pretraining corpora and similar training process Xia et al. (2024).
To clarify the above statements, we compare our method with EAGLE-2 Li et al. (2024b), a typical tiny-drafter speculative decoding method. EAGLE-2 uses a well-trained one-layer draft model (the parameter size is around 1B), aiming for extreme drafting latency reduction. We evaluated both methods on Spec-Bench Xia et al. (2024), where EAGLE-2 outperforms all existing methods according to the leaderboard. Besides average token throughput, we also calculate the variances of throughput across data items, for better demonstration of stability and generalization capability. The temperature is set to 0 for both methods.
Firstly we choose Llama-3-70B-Instruct as the base model. The results in the top part of Table 4 show that, the EasySpec-accelerated 8B drafter achieves throughput comparable to, or even surpassing, the well-trained drafter of EAGLE-2, with no training or fine-tuning on either the target distribution or fuzzy approximation. This suggests that a smaller model from the same series, if existing, is good enough to deliver satisfactory performance. On the other hand, EasySpec exhibits significantly smaller variances across all datasets, indicating that a relatively larger draft model can offer more stable and predictable performance. We attribute this stability and generalization capability to the higher speculation accuracy, which reaches nearly 0.8 in contrast to the less than 0.4 accuracy of EAGLE-2. Even trained upon the target tasks, the smaller drafter is inherently incapable of highly accurate speculation due to its limited parameter size.
To further demonstrate the generalization capability across base models, we tested both methods on Llama-3.3-70B-Instruct, as there is no smaller model in its series and both methods need to generalize. The results in the bottom part of Table 4 illustrate that EasySpec greatly outperforms EAGLE-2 across all types of tasks, with less than a 10% drop of accuracy compared to the upper results. In contrast, both the average throughput and the acceptance rates of EAGLE-2 drop significantly.
Table 4: Average token throughput (variances) and acceptance rates of EasySpec and EAGLE-2 on Spec-Bench.
EAGLE-2 EasySpec
Task Type Throughput $\alpha$ Throughput $\alpha$
Llama-3-70B-Instruct
MT 33.81(27.09) 0.35 31.33(4.75) 0.80
TL 30.44(13.12) 0.34 31.69(6.86) 0.83
SUM 33.26(11.30) 0.36 31.13(3.34) 0.83
QA 30.98(21.69) 0.31 31.32(6.80) 0.79
MR 34.37(12.84) 0.39 34.43(4.72) 0.85
RAG 33.7(32.99) 0.40 29.91(8.19) 0.83
Llama-3.3-70B-Instruct
MT 23.38(13.86) 0.20 29.52(15.50) 0.76
TL 18.60(5.96) 0.14 28.00(12.43) 0.73
SUM 23.22(4.16) 0.19 31.05(4.48) 0.83
QA 21.31(6.35) 0.17 29.29(9.46) 0.75
MR 26.95(5.78) 0.25 33.79(6.36) 0.84
RAG 23.03(16.41) 0.20 28.45(16.65) 0.77
Figure 4: Token throughput and acceptance rates of Llama-3-70B(8B)-Instruct on MMLU under different configurations of EasySpec.
4.3 Approximation Precision
The feasibility of fuzzy speculation relies on the approximation precision of hidden states inside the draft model, which can be measured by cosine similarity. Table 5 shows the average cosine similarities between the precise and fuzzy-approximated hidden states. The tested drafter is Llama-3-8B-Instruct and the benchmark is MMLU. The evaluated hidden states include input hidden states $h$, outputs of attention layers $Attnoutput$, and queries(q), keys(k) and values(v) inside attention heads. We set layer-parallel sizes to 2,3,4 for diverse approximation granularity. The results show that, as more layers are parallelized, the cosine similarity decreases but stays above 0.8, indicating a sufficiently high precision of approximation. Notably, the queries and keys are well-approximated with cosine similarities approaching 1. This is critical for maintaining precision of other hidden states and outputs, as errors in attention weights scale exponentially.
Table 5: Average cosine similarities between precise and approximated hidden states. The tested model and testbench are Llama-3-8B-Instruct and MMLU.
LP size $h$ q k v $Attnoutput$
2 0.93 0.98 0.99 0.92 0.93
3 0.89 0.97 0.98 0.86 0.88
4 0.86 0.96 0.97 0.82 0.83
4.4 Ablation study
We conducted an ablation study on 3 aspects of EasySpec: tree attention width, layer-parallel (LP) size and the presence or absence of calibration. We varied the tree width across 1,4,8,12 and LP size from 1 to 5. We measure token throughput and acceptance rates of Llama-3-70B(8B)-Instruct on MMLU under these different configurations, and present the results in Figure 4.
As the LP size increases from 1, the token throughput improves due to better multi-GPU utilization of fuzzy speculation. However, when the LP size becomes too large, the speculation results are excessively fuzzy, leading to a noticeable decline in acceptance rates and hence token throughput. The optimal LP size is influenced by the tree width: a wider tree increases the likelihood of existence of correct tokens, thus allowing for a fuzzier and faster speculation (a larger LP size). On the other hand, even without tree attention (width=1), EasySpec can also accelerate the inference with layer parallelism (see the green line in Figure 4).
Bonus calibration plays a critical role in preserving speculation accuracy. When the LP size is 1, calibration is unnecessary, since there are no approximation errors in the KV cache. However, as the LP size increases, for every combination of tree width and LP size, bonus calibration can significantly improve token throughput and acceptance rates. Bonus calibration also slows the decline in acceptance rates with increasing LP sizes (see the right part of Figure 4). All the above results prove that the trade-off between precision of KV values and additional latency is reasonable and effective.
5 Related Work
Existing works on speculative decoding mainly focus on the drafting stage. Methods include training efficient draft models, self-speculation or algorithmic approaches. Specinfer Miao et al. (2024) uses boost-trained draft models and tree-structured attention for efficient speculation. Medusa Cai et al. (2024) trains a set of extra MLP heads for future token prediction using the original LLM’s features. Self-Speculative Decoding Zhang et al. (2023) and LayerSkip Elhoushi et al. (2024) skips some layers of the target LLM for self-speculation. Lookahead Fu et al. uses n-gram Jacobi decoding to increase acceptance rates. REST He et al. (2023) leverages a data storage for retrieval speculation.
The discrepancy between optimal distribution sizes is discovered by the authors of Chen et al. (2023), who suggested training a wider but shallower draft model for better performance under tensor parallelism. Few works focus on this problem, yet it is important and worth-solving. To the best of our knowledge, we are the first to effectively optimize multi-GPU utilization of currently-available draft models.
6 Conclusion
In this paper, we propose EasySpec, a layer-parallel speculation strategy designed to optimize the efficiency of multi-GPU utilization during the drafting stage of speculative decoding. EasySpec introduces two modifications to the original speculation process: fuzzy speculation and bonus calibration. Fuzzy speculation breaks the sequential layer execution order and enables multi-layer parallelization, while bonus calibration applies token parallelism to the drafting stage for mitigating long-term error accumulation. EasySpec achieves consistent acceleration across models and datasets, requiring no training or fine-tuning on the existing draft models.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Speculative Decoding
Speculative decoding consists of iterations of drafting and verification. The detailed algorithm of a single iteration is present in Algorithm 3.
Algorithm 3 An Iteration of Speculative Decoding
  Input: input token sequence $X$, drafting length $m$, drafting model $M^{\prime}$, base model $M$
  for $i=1$ to $m$ do
     $x^{\prime}_{i},p^{\prime}_{i}=M^{\prime}(X,x^{\prime}_{j<i})$
  end for
  $p_{1},\cdots,p_{m},p_{m+1}=M(X,[x^{\prime}_{1},\cdots,x^{\prime}_{m}])$
  sample $r_{1},\cdots,r_{m}$ independently from $U(0,1)$
  for $i=1$ to $m$ do
     if $r_{i}>\frac{p_{i}(x^{\prime}_{i})}{p^{\prime}_{i}(x^{\prime}_{i})}$ then
        $n=i-1$ ($n$ is the accepted length)
        break (rejected)
     else
        $x_{i}=x^{\prime}_{i}$ (accepted)
     end if
  end for
  if $n<m$ then
     $p_{bonus}=norm(\max(0,p_{n}-p^{\prime}_{n}))$
  else
     $p_{bonus}=p_{n+1}$
  end if
  sample $x_{bonus}$ from $p_{bonus}$
  return $X,x_{1},\cdots,x_{n},x_{bonus}$
For any distribution $p^{\prime}$, the final output token sequence is equivalent to sampling them directly from $p$. The proof is specified in many existing works (e.g. Leviathan et al. (2023)).
Appendix B Layer-Parallel Strategies
The detailed layer-parallel strategies of different models and LP sizes are listed in Table 6. The ’a-b’ in the table represents that the input hidden states $h_{a},h_{a+1},\cdots,h_{b}$ are all equal to $h_{a}$. We empirically discovered that the input of layer 0 (the word embedding) can hardly approximate input of other layers, so layer 0 is never bounded with any other layers. All the tested draft models from the Qwen family have 28 layers, and the parallel strategy for them is identical.
Table 6: Layer-parallel strategies for draft models. ’Qwen’represents all Qwen draft models.
Model $N$ Strategy
Llama-3-8B-Instruct 2 0, 1-2, 3-4, …, 29-30, 31
3 0, 1-3, 4-6, …, 25-27, 28-30, 31
4 0, 1-3, 4-7, 8-11, …, 24-27, 28-30, 31
5 0, 1-3, 4-8, 9-13, …, 24-28, 29-30, 31
Qwen 4 0, 1-3, 4-7, 8-11, …, 24-26, 27
Note that the optimal layer-parallel strategy depends on specific models and datasets, and we have not elaborated on finding more fine-grained strategies. Modifying these strategies could lead to improved performance.

### v2

@@ Appendix G Broader impacts
EasySpec 1.61($\uparrow$1.57x) 1.63 3.24(3.64x) 0.87 1.48($\uparrow$1.52x) 1.54 3.02(3.97x) 0.91
@@
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The paper provides EasySpec, a layer-parallel speculation strategy for efficienct multi-GPU utilization during the drafting stage of speculative decoding, introducing two modifications to the original speculation process: fuzzy speculation and bonus calibration. The abstract and introduction clearly state these contributions
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The paper discusses the limitation of the work in Appendix˜F, where the limitation of specific software and hardware configuration is discussed.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: There is no theoretical result in the paper. We did provide theoretical backing for our method in Appendix˜E, while it is stated for further clarification of the empirical results. Therefore, we do not think it is in any way a theoretical result.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We believe that the detailed experimental settings are provided, in Section˜4 and some places in the paper. They include model types, datasets, layer parallel strategies, speculation length, and so on.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: The code and the benchmarks are all available in the repository link.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: As we mentioned in ’Experimental result reproducibility’, we believe that detailed hyperparameters have been provided for understanding.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: Similar to many existing works in this area, we do not include error bars or confidence intervals in the empirical results, as repeated trials show negligible variation. System-level fluctuations have only marginal effects on the outcomes.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: The system configuration is specified in Section˜4 and Appendix˜F..
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The research conducted in the paper conform with the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: The broader impacts are discussed in Appendix˜G.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: To the best of our knowledge, this paper poses no issues on safeguards.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We cited original papers for code, datasets or content. The license and terms of use are explicitly mentioned and properly respected.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We provide assets in the form of code, which is well documented and commented alongside.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The core method development in this research does not involve LLMs as any important, original, or non-standard components.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.